# Learning a Distributed Hierarchical Locomotion Controller for Embodied Cooperation

### v1

@@ Abstract
In this work, we propose a distributed hierarchical locomotion control strategy for whole-body cooperation and demonstrate the potential for migration into large numbers of agents. Our method utilizes a hierarchical structure to break down complex tasks into smaller, manageable sub-tasks. By incorporating spatiotemporal continuity features, we establish the sequential logic necessary for causal inference and cooperative behaviour in sequential tasks, thereby facilitating efficient and coordinated control strategies. Through training within this framework, we demonstrate enhanced adaptability and cooperation, leading to superior performance in task completion compared to the original methods. Moreover, we construct a set of environments as the benchmark for embodied cooperation. A visual depiction of highlights of the learned cooperative behaviours can be found on this website [^1] Videos of our results are available at our project webpage: https://d-hrl.github.io.
@@ 1 Introduction
Cooperatively accomplishing embodied tasks by multiple robots has consistently been a highly challenging area of research. Recent studies mainly focus on embodied manipulation cooperation among robotic arms or formation control over the upper level within a group of mobile robots [1, 2]. Nevertheless, multi-agent cooperation via whole-body and end-to-end locomotion control is rarely studied.
Some previous works showcase the manipulation via locomotion [3] but are only tested on two agent systems, and the scalability of this method is still agnostic for migration to any number of agent populations. In this work, we aim to realize more complex embodied multi-agent cooperation by learning a distributed hierarchical locomotion control system, decomposing the complex and coupled behaviours while maintaining the potential for unlimited expansion on the swarm. As the foundation for implementation and validation, we construct three scenarios in IsaacSim/Gym [4] as benchmarks for embodied cooperation study.
@@
Concurrently, training a robot for a specific function can be effectively achieved through reinforcement learning (RL), like learning movement patterns [5], interactive behaviours [6], as well as logical inference in games [7]. Although RL provides a recognized powerful exploration capability and tremendous progress has been made in sampling efficiency [4], finding and mastering a sequence of sophisticated tasks through searching remains a challenging problem. Hierarchical reinforcement learning (HRL) alleviates this to a certain extent, aiming to understand the logical relationships among ”control, action, behaviour, dynamic outcomes, and feedback” in a segmented manner. An increasing number of studies are employing hierarchical approaches to decompose complex tasks [8, 9, 10, 11, 12, 13, 14, 15]. Inspired by these, we propose a three-level HRL to naturally decompose perception feature extraction, cooperative behaviours, and fundamental actions.
In the previous work [3], embodied cooperation is completed via locomotion control by two quadruped robots. However, spatiotemporal features between a sequence of events are not considered in this work, which causes failure in cooperation of more than two agents. Actually, understanding interactive and cooperative behaviours requires sequential causal reasoning considering world dynamics. To enhance this, we deploy sequential behaviour learning with spatiotemporal memory capabilities in the middle layer of the hierarchical network by constructing a Recurrent Neural Network (RNN). Furthermore, the previous method can hardly be applied to a flock of agents [3], while the scalability of the swarm is crucial in multi-agent reinforcement learning (MARL) field [16, 17]. Herein, we construct our method facing the decentralized and distributed scenario and attempt to apply it to an expandable MAS.
@@ 3 Problem Statement
We first formulate the problem of learning embodied cooperative behaviour in a swarm of agents, then try to solve this problem with MARL. Since the number of agents is fixed within an episode, the synchronous framework and policy gradient descent algorithms are selected. To realize real-time execution, we further select the Centralized Training and Decentralized Execution (CTDE) framework to support our implementation. It is crucial to mention that agents in the group might be homogeneous or heterogeneous, and agents in the same species share the same policy.
@@ 4 Distributed Hierarchical Reinforcement Learning
As aforementioned, locomotion control for embodied cooperation is an extremely challenging task. There are three main difficulties: (i) How to decompose and decouple the complicated tasks into a step-by-step mode? (ii) How to establish the temporal and spatial correlation of behaviours? (iii) Whether the trained models can expand to a large scale of multiagent collaboration? In this part, we will address each of these three questions in detail.
@@ 4.2 Spatiotemporal Memory Recurrence
Capturing and leveraging spatiotemporal features is significant for long-term decision-making tasks [40]. For tasks requiring cooperation, spatial awareness is as crucial as understanding temporal dynamics. We propose spatiotemporal memory recurrence that uses RNNs to enhance the agent’s ability to make informed decisions over extended periods and complex environments.
The significance of learning spatiotemporal features through RNNs in distributed HRL lies in the inherent capability of these networks to maintain and utilize historical information. This capability is particularly important in dynamic environments, especially in nonstationary multiagent environments. RNNs, on the other hand, are designed to handle sequential data, making them ideal for tasks that require an understanding of the temporal progression of states and actions. Thus, we maintain a hidden recurrent state $h$ to pass the feature of scenarios continuously, as shown in Fig. 2.
4.3 Distributed Scalable HRL
In contrast to previous studies that employed HRL or end-to-end approaches for robotic cooperative manipulation or locomotion, our distributed HRL method considers the capability of scaling the population. To achieve this, we need to strictly define the format of input and output information for each agent, particularly the observation. Herein, we propose to use partial observation of agents related to their colleagues, according to their visibility [41]. Each agent can only obtain the nearest neighbour’s local position. As illustrated in Fig.2, the dimension of exteroceptive state $e$ is limited. This actually changes the difficulty of collaboration learning. Since partially observed agents can only see nearby peers and environmental information, they need to estimate the states of other invisible agents from the dynamics of the environment and predict future changes in the environment.
@@ 4.4 Training Curriculum
Acquiring collaborative behaviours requires the agents’ random search to span a sufficiently large space, which in complex tasks can often lead to an inability to learn usable knowledge due to the vastness of the search space. In this case, we design some appropriate training curricula that reasonably use sparse rewards and dense rewards in the early stages to intentionally guide the agent to approach success. This curriculum is independent of HRL pertaining, such as encouraging agents to move to the target first at the early stage. The curriculum designs are described in detail in the appendix.
@@ 5.1 Environments Construction
Three scenarios are established that represent different task types: Cooperative Transport, Corridor Crossing, and Ravine Bridging.
@@ 5.2.1 Training Basic Locomotion Control
In the training of lower-level motion control, two control paradigms are available: velocity control and position control. We first train the two control paradigms in the single-agent environment to obtain accurate control of the fundamental actions. We give a reciprocal $L_{2}$ distance between the current position and the target position, or the dot product of the agent velocity and target velocity, as the main reward in each training curriculum. Fig 4(a) shows the training results of two paradigms. We find that the velocity control offers greater flexibility than position control. In our empirical tests, however, position control presents higher accuracy. Moreover, velocity control makes it difficult to maintain a stationary position at a specific point. Similar to any other hierarchical learning study, the lower layer is frozen once its training is done well.
@@ 5.2.2 Training the Interactive Collaboration
After the lower layer is elaborately trained, we freeze this network and fix the basic locomotion control performance. Then, the upper and middle layers are trained on different scenes via distinct reward mechanisms. The training results are shown in Fig. 4(b,c,d).
@@
Ravine Bridging: The group in this task is a heterogeneous team. Although all agents are Ant, there are two sets of policies: one for agents on the plane and one for agents in the ravine. Agents on the plane will arrive at the edge of ravine and stop, waiting for the agents in the ravine to transport the bridge to the front of them. We get approximately half success over one hundred trails. The main reason is that Ants in the ravine might get stuck or fall.
Table 1: Success Rate Comparison: We present success rates with standard error on the three tasks, comparing our method to ablations. Each percentage results from 50 trials—10 attempts per three independently trained models.
@@
Cooperative Transport $\textbf{74}\pm 7\%$ $0\pm 0\%$ $5\pm 1\%$
Corridor Crossing $\textbf{93}\pm 3\%$ $0\pm 0\%$ $2\pm 1\%$
Ravine Bridging $\textbf{48}\pm 5\%$ $0\pm 0\%$ $0\pm 0\%$
Figure 5: Illustration of hierarchical learning concepts applied in different scenarios compared with no hierarchy case, noisy hidden states case and the baseline [3] which does not have the middle layer with recurrent neural networks. All results are conducted on four agents scenario.
@@ 5.3.1 The role of Hierarchical Learning
In this experiment, we replace the hierarchical networks with a single-layer MLP, shown in Fig. 5. We find the reward converges to an extremely low level quickly. Agents can only learn and get the basic reward of locomotion. These results show that hierarchy is crucial to success in more complex tasks that require coordination. Without hierarchy, agents cannot complete the mission (in Table. 1). The reason for failure is obvious. As with other HRL objectives, a single network facing high-dimensional observations and complex multi-item reward functions finds it difficult to distinguish features between channels and learn behaviour guided by a single main reward.
For multi-agent training, the situation is even more unique. Due to the CTDE training paradigm we default to, differences in the positions of multiple bodies lead to an inability to uniformly learn certain distance rewards. For example, in the task of cooperative transport, agents have different positions compared to objects and targets, so a unified strategy cannot be directly formed by lower-level networks. It can be said that hierarchical learning is indispensable for such group tasks.
@@ 5.3.2 The effectiveness of Spatiotemporal Memory
To validate the effectiveness of spatiotemporal features, we re-product the method in previous work [3] as the baseline, which does not consider the event correlation. As shown in Fig.5, our method obtains a much higher reward for all three tasks. The approach proposed in  [3] which does not have RNNs, converges to a slightly higher reward than in the case of no hierarchy but is still far from completing the mission. Furthermore, we attempt to replace the RNN hidden states with random noise inputs to explain how event correlation positively and effectively affects the entire performance. The success rate of all these attempts is close to zero. We find that agents cannot learn a set of systematic behaviours in specific tasks, and there are occasional successes when the object position is really close to the target. Actually, if the agents only react to the current observation, they could never form strategic behaviours. In other words, he would never think for himself in the next second without spatiotemporal memory. This is even more deadly in collaborative tasks, as we can obviously find that the conciliatory and sacrificial behaviour emerging in crossing corridor tasks is a strategy of thinking for long-term consequences.
5.4 Analysis
Scalability Analysis: We validate scalability in the task cooperative transport as illustrated in Fig. 6(a). The final reward increases as the number of agents increases, which also aligns with our expectations for this task. More agents participating should indeed bring better transportation results. Our results reveal that the distributed HRL has great potential to migrate a trained policy to a large number of agents. Only such algorithms can achieve the true sense of ”biocoenosis” collaborative tasks.
@@ 5.4 Analysis
(b) Distributed HRL
(c) No hierarchy
(d) No memory
Figure 6: (a) shows the training results of different Ant populations used in Cooperative Transport. (b)(c)(d) shows the success and failure distribution of task Cooperative Transport.
Typical Result Analysis: We analyse the typical result of task cooperative transport. To test the robustness of our method, we conduct more than one hundred trials with randomized reset positions and count the success rate then plot the scatter diagram, as illustrated in Fig. 6(b). Each point represents an initialized position of an agent, and the object is located at (0,0) point. Red points denote agents generated at these positions fail to complete the task, while each blue point denotes a success. We can observe that when the initial position is closer to the object, the success rate tends to be higher, and there is only one failure case within a 5-meter range of the object; conversely, the success rate decreases as the distance from the object increases. This is because the probability of accidents is greatly reduced by short-distance tasks. It is worth mentioning that most of the failure cases result from the agent tipping over or pushing the object obviously to the side with uneven force during the pushing process. Fig.6(c)(d) showcase the evaluation results under no-hierarchy and no-spatiotemporal memory. Only few successes can be found in the evaluation of no-spatiotemporal memory, and all cases fail in training without hierarchy.
6 Conclusion
In this work, we propose a distributed HRL architecture for embodied cooperation, and the results showcase the effectiveness of our method in embodied and whole-body control collaborative tasks. Besides, our approach has significant potential in scaling up the flock size. In the future, more complex cooperative tasks can be conducted based on our method.
@@ Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
@@ A.1.1 Cooperative Transport
The agents should cooperate to push a cylinder to a target location. The local positions between the agents, cylinder, and target are initialized randomly. The cylinder is set in the middle of the map, and agents are initialized in the four quadrants of the cylinder respectively.
@@ A.1.3 Ravine Bridging
A.2 Curriculum
We utilize sparse rewards to direct agents towards proximity with the object, subsequently acquiring the target’s location in the Cooperative Transport task. In the Corridor Crossing task, we train individual agents to navigate to the finish line by passing through the slit and then refine the pre-trained policy for application in a multi-agent environment. The learning sequence for the Ravine Bridging task is more intricate which means we have to give more guidance; we instruct agents to wait at the ravine’s edge at the opportune moment to prevent falling, while other agents within the ravine initially learn to convey ”bridge” to an appropriate location.
A.3 Scalability
The distributed HRL framework allows the collaborative tasks to be explicitly represented as a ”group”, thereby enabling tasks to be extended to large numbers of agents team. We have tested the framework with different populations of Ant in various tasks (There are only two Ants for pushing bridge in the task Ravine Bridging), and the success rates of each task with different populations are shown in the Tab. 2. The results show the potential of our framework for handling tasks with more agents.
Table 2: Each percentage results from 50 trials—10 attempts per three independently trained models.
Success rate Two Ants Three Ants Four Ants Five Ants
Cooperative Transport $8\pm 2\%$ $42\pm 3\%$ $74\pm 7\%$ $82\pm 5\%$
Corridor Crossing $98\pm 2\%$ $95\pm 2\%$ $93\pm 3\%$ $79\pm 2\%$
Ravine Bridging $48\pm 5\%$ $36\pm 8\%$ $30\pm 2\%$ $12\pm 5\%$
Table 3: Success Rate Comparison between average moving distances in task Ravine Bridging.
Distance from Ravine Two Ants Three Ants Four Ants Five Ants
2 m $49\%$ $45\%$ $38\%$ $35\%$
4 m $38\%$ $33\%$ $20\%$ $11\%$
6 m $30\%$ $19\%$ $5\%$ $3\%$
A.4 Typical Result
A.4.1 Corridor Crossing

### v2

@@ Abstract
In this work, we propose a distributed hierarchical locomotion control strategy for whole-body cooperation and demonstrate the potential for scaling to large numbers of agents. Our method utilizes a hierarchical structure to break down complex tasks into smaller, manageable sub-tasks. By incorporating spatiotemporal continuity features, we establish the sequential logic necessary for causal inference and cooperative behaviour in sequential tasks, thereby facilitating efficient and coordinated control strategies. Through training within this framework, we demonstrate enhanced adaptability and cooperation, leading to superior performance in task completion compared to the original methods. Moreover, we construct a set of environments as the benchmark for embodied cooperation. A visual depiction of highlights of the learned cooperative behaviours can be found on this website [^1] Videos of our results are available at our project webpage: https://d-hrl.github.io.
@@ 1 Introduction
Cooperatively accomplishing embodied tasks with multiple robots has consistently been a highly challenging area of research. Recent studies mainly focus on embodied manipulation cooperation among robotic arms or formation control over the upper level within a group of mobile robots [1, 2]. Nevertheless, multi-agent cooperation via whole-body and end-to-end locomotion control is rarely studied.
Some previous works showcase the manipulation via locomotion [3] but are only tested on two agent systems, and the scalability of this method is still agnostic for migration to any number of agent populations. In this work, we aim to realize more complex embodied multi-agent cooperation by learning a distributed hierarchical locomotion control system, decomposing the complex and coupled behaviours while maintaining the potential for unlimited expansion on the swarm. As the foundation for implementation and validation, we construct three scenarios in IsaacSim [4] as benchmarks for embodied cooperation study.
@@
Concurrently, training a robot for a specific function can be effectively achieved through reinforcement learning (RL), such as learning movement patterns [5], interactive behaviours [6], as well as logical inference in games [7]. Although RL provides a recognized powerful exploration capability and tremendous progress has been made in sampling efficiency [4], finding and mastering a sequence of sophisticated tasks through searching remains a challenging problem. Hierarchical reinforcement learning (HRL) alleviates this to a certain extent, aiming to understand the logical relationships among ”control, action, behaviour, dynamic outcomes, and feedback” in a segmented manner. An increasing number of studies are employing hierarchical approaches to decompose complex tasks [8, 9, 10, 11, 12, 13, 14, 15]. Inspired by these, we propose a fully-decentralized HRL framework to naturally decompose perception feature extraction, cooperative behaviours, and fundamental actions.
In the previous work [3], embodied cooperation is completed via locomotion control by a small number of robots. However, as the number of agents increases, previous methods fail to effectively address the issue, while the scalability of the swarm is crucial in multi-agent reinforcement learning (MARL) field [16, 17, 18, 19]. To tackle this challenge, we propose a fully-decentralized framework that can be applied to a flock of agents to enhance the method’s scalability and promote impressive cooperative behavior among larger-scale intelligent agents. We aim to apply this framework to an expandable MAS. Moreover, spatiotemporal features between a sequence of events are not considered in the previous work, which causes failure in understanding interactive and cooperative behaviours. To enhance this, we deploy sequential behaviour learning with spatiotemporal memory capabilities in the middle layer of the hierarchical network by constructing a Recurrent Neural Network.
@@ 3 Problem Statement
We first formulate the problem of learning embodied cooperative behaviour in a swarm of agents, then try to solve this problem with MARL. Since the number of agents is fixed within an episode, the synchronous framework and policy gradient descent algorithms are selected. To realize real-time execution, we further proposed a fully-decentralized training framework to support our implementation. It is crucial to mention that agents in the group might be homogeneous or heterogeneous, and agents in the same species share the same policy.
@@ 4 Distributed Hierarchical Reinforcement Learning
As aforementioned, locomotion control for embodied cooperation is an challenging task. There are three main difficulties: (i) How to decompose and decouple the complicated tasks into a step-by-step mode? (ii) Whether the trained models can expand to a large scale of multiagent collaboration? (iii) How to establish the temporal and spatial correlation among collective behaviors? In this part, we will address each of these three questions in detail.
@@ 4.1 Task Decomposition and Hierarchical Learning
4.2 Distributed Scalable HRL
In contrast to previous studies that employed HRL or end-to-end approaches for robotic cooperative manipulation or locomotion, our distributed HRL method considers the capability of scaling the population. To achieve this, we need to strictly define the format of input and output information for each agent, particularly the observation. Herein, we propose to use partial observation of agents related to their colleagues, according to their visibility [42]. Each agent can only obtain the nearest neighbour’s relative position. As illustrated in Fig.2, the dimension of exteroceptive state $e$ is limited. This actually changes the difficulty of collaboration learning. Since partially observed agents can only see nearby peers and environmental information, they need to estimate the states of other invisible agents from the dynamics of the environment and predict future changes in the environment. However, such observation settings theoretically enable our framework to scale to a significantly large amount of agents system, which is unattainable by CTDE frameworks.
@@ 4.3 Spatiotemporal Memory Recurrence
Capturing and leveraging spatiotemporal features is crucial for long-term decision-making tasks [43] particularly the complex embodied cooperation tasks we designed. We anticipate that fully-decentralized agents should possess the capability to comprehend their own and other agents’ continuous behaviors, thus forming collaborative relationships in space across continuous time intervals. Furthermore, this capability holds particular significance in dynamic environments, especially within fully-decentralized nonstationary multi-agent environments. Therefore, we propose spatiotemporal memory recurrence utilizing RNNs to enhance the agent’s ability to make informed decisions over extended periods and complex environments. Consequently, we maintain a hidden recurrent state $h$ to pass the feature of scenarios continuously, as shown in Fig. 2.
@@ 4.4 Training Curriculum
Acquiring collaborative behaviours requires the agents’ random search to span a sufficiently large space, which in complex tasks can often lead to an inability to learn usable knowledge due to the vastness of the search space. In this case, we design some appropriate training curricula that reasonably use sparse rewards and dense rewards in the early stages to intentionally guide the agent to approach success. The curriculum designs are described in detail in the appendix.
@@ 5.1 Environments Construction
Three scenarios are established that represent different task types: Cooperative Transport, Corridor Crossing and Ravine Bridging.
@@ 5.2.1 Training Basic Locomotion Control
In the training of lower-level motion control, two control paradigms are available: velocity control and position control. We first train the two control paradigms in the single-agent environment to obtain accurate control of the fundamental actions. We give a reciprocal $L_{2}$ distance between the current position and the target position, or the dot product of the agent velocity and target velocity, as the main reward in each training curriculum. Fig 4(a) shows the training results of two paradigms. After training, We find that the velocity control offers greater flexibility than position control. However, position control presents higher accuracy. Moreover, velocity control makes it difficult to maintain a stationary position at a specific point.
@@ 5.2.2 Training the Interactive Collaboration
After the lower layer is elaborately trained,we adopt the position control paradigm and freeze this network. Then, the upper and middle layers are trained on different scenes via distinct reward mechanisms. The training results are shown in Fig. 4(b,c,d).
@@
[…]
Table 1: Success Rate Comparison: We present success rates with standard error on the three tasks, comparing our method to ablations. Each percentage results from 500 trials — 100 attempts per five independently trained models.
@@
Cooperative Transport $\textbf{76.8}\pm\textbf{5.3}\%$ $0.0\pm 0.0\%$ $9.6\pm 3.2\%$
Corridor Crossing $\textbf{88.4}\pm\textbf{4.8}\%$ $0.0\pm 0.0\%$ $12.0\pm 4.1\%$
Ravine Bridging $\textbf{50.4}\pm\textbf{4.4}\%$ $0.0\pm 0.0\%$ $5.4\pm 4.2\%$
@@ 5.3.1 The role of Hierarchical Learning
In this experiment, we replace the hierarchical networks with a single-layer MLP. Given no hierarchical structure, there is no form of target command and velocity command, which precludes the possibility of training an identical lower-level controller in advance. Therefore, when training without hierarchy, we provide the same training curriculum tailored to each task, as comprehensively detailed in the Appendix. Furthermore, we provide proprioceptive states and exteroceptive states as observations and the same task reward and locomotion reward as the total rewards.
We find that agents can only learn simple strategy and get the basic reward which are easy to reach. These results show that hierarchy is crucial to success in more complex tasks that require coordination. Without hierarchy, agents cannot complete the mission (in Table. 1). The reason for failure is obvious. As with other HRL objectives, a single network facing high-dimensional observations and complex multi-item reward functions finds it difficult to distinguish features between channels and learn behaviour guided by a single main reward. Especially, we need a frozen lower-level locomotion controller, which is the foundation for success.
@@ 5.3.2 The effectiveness of Spatiotemporal Memory
To validate the effectiveness of spatiotemporal features, we re-product the method in previous work [3] as the baseline, which does not consider the event correlation. We provide the same low-level locomotion controller. The approach proposed in  [3] which does not have RNNs, converges to a slightly higher reward than in the case of no hierarchy but is still far from completing the mission. As shown in Table. 1, the success rate of this method maintains a relatively low level which rarely exceeding 15%. We find that agents cannot learn a set of systematic behaviours in specific tasks, and there are occasional successes. Actually, if the agents only react to the current observation, they could never form strategic behaviours. In other words, he would never think for himself in the next second without spatiotemporal memory. This is even more deadly in collaborative tasks, as we can obviously find that the conciliatory and sacrificial behaviour emerging in crossing corridor tasks is a strategy of thinking for long-term consequences.
5.4 Scalability Experiments
We emphasize the advantages of our proposed fully-decentralized approach in terms of scalability, which includes the advantage of training when the number of agents is large, as we do not need to provide the global observations of all agents like CTDE framework during training to avoid the huge increase in training complexity. On the other hand, our framework also has advantages at the ability of directly transferring the trained model to more agents tasks.
First, we demonstrate the capability of our framework to effectively train agents with nice performance as the number of agents increases, even when these agents only have access to partial observations. We validate this in the task cooperative transport as illustrated in Fig. 5(a). The final reward increments as the number of agents rises, which is also in line with our expectations for this task. The participation of more agents should indeed lead to better transportation results.
Second, we conducted zero-shot transfer tests on our trained model (trained with four agents) across diverse populations of Ant in each task(Agent number indicates the agent on the plane in the task Ravine Bridging), and the success rates of each task with different populations are shown in the Tab. 5.4. The outcomes demonstrate the potential of our framework’s transfer capability.
Success rate Two Ants Three Ants Four Ants Five Ants
Cooperative Transport $12.8\pm 5.6\%$ $39.8\pm 4.5\%$ $\textbf{76.8}\pm\textbf{5.3}\%$ $82.2\pm 4.5\%$
Corridor Crossing $97.1\pm 1.8\%$ $93.6\pm 2.2\%$ $\textbf{88.4}\pm\textbf{4.8}\%$ $73.5\pm 2.5\%$
Ravine Bridging $\textbf{50.4}\pm\textbf{4.4}\%$ $36.2\pm 7.8\%$ $22.3\pm 5.1\%$ $10.9\pm 3.3\%$
Third, we conduct the experiment to show the advantages of our fully decentralized method training advantage over the CTDE methods when the number of agents increases. The most appropriate task to show is Corridor Crossing because as the number of agents increases, the difficulty of the task Cooperative Transport decreases, while Ravine Bridging is too complex and the completion rate is too low when the number of agents is large. So we add the number of agents to 6 and 7 in corridor crossing. And the Fig. 5(b)(c) shows that the reward of centralized training grows slower and ultimately not converge.
@@ 5.4 Scalability Experiments
(b) Training with six agent number
(c) Training with seven agent number
Figure 5: (a) shows the training results of different Ant populations used in Cooperative Transport. (b)(c)shows the result training with large agent numbers.
6 Conclusion and Limitations
In this work, we propose a distributed HRL architecture for embodied cooperation, and the results showcase the effectiveness of our method in embodied and whole-body control collaborative tasks.
Despite the encouraging findings, it is essential to recognize the limitations of our research. Firstly, although the hierarchical network method endows our framework with a certain ability to address long-horizon problems, its capacity to solve complex issues remains limited as all layers are still fully trained based on reinforcement learning. We primarily relied on continuous reward shaping and increasing the total training steps to obtain better training outcomes. We contend that if we aim to solve more complex long-horizon cooperative tasks with the participation of more intelligent agents, additional planning-based guidance might be necessary at the upper level.
Moreover, our work is indeed more algorithmically and simulation benchmark-oriented, and we have not explored robots with more complex morphology. Despite having already demonstrated the promising results of our method on some tasks and witnessed the emergence of cooperative behavior, there is still a considerable engineering workload required in terms of control and sim2real transfer.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grants 62025304.
@@ A.1.1 Cooperative Transport
The agents should cooperate to push a cylinder to a target location. The target position are initialized randomly. The cylinder is set in the middle of the map, and agents are initialized in the four quadrants of the cylinder respectively.
@@ A.1.3 Ravine Bridging
A.2 Typical Result
A.2.1 Cooperative Transport
We analyse the typical result of task cooperative transport. To test the robustness of our method, we conduct more than one hundred trials with randomized reset positions and count the success rate then plot the scatter diagram, as illustrated in Fig. 6(b). Each point represents an initialized position of an agent, and the object is located at (0,0) point. Red points denote agents generated at these positions fail to complete the task, while each blue point denotes a success. We can observe that when the initial position is closer to the object, the success rate tends to be higher, and there is only one failure case within a 5-meter range of the object; conversely, the success rate decreases as the distance from the object increases. This is because the probability of accidents is greatly reduced by short-distance tasks. It is worth mentioning that most of the failure cases result from the agent tipping over or pushing the object obviously to the side with uneven force during the pushing process. Fig.6(c)(d) showcase the evaluation results under no-hierarchy and no-spatiotemporal memory. Only few successes can be found in the evaluation of no-spatiotemporal memory, and all cases fail in training without hierarchy.
(a) Distributed HRL
(b) No hierarchy
(c) No memory
Figure 6: (a)(b)(c) shows the success and failure distribution of task Cooperative Transport.
A.2.2 Corridor Crossing